OpenStudy (anonymous):

One leg of the right triangle is 20 centimeter and the hypotenuse is 10 centimeter longer than the other leg. Find the lengths of the unknown sides.

OpenStudy (anonymous):

pythagoras for this one put the length of the "other side" as x, then hypotenuse is \[x+10\] two legs are 20 and x pythagoras says \[20^2+x^2=(10+x)^2\] and now you have a nice equation to solve

OpenStudy (anonymous):

how?? please give me the exact soultion

OpenStudy (anonymous):

it looks like a quadratic equation but it is not. square and get \[400+x^2=100+20x+x^2\] i bet you can go from there right?

OpenStudy (anonymous):

it's not hard just a lot of leg work

OpenStudy (anonymous):

welcome to "real" math...enjoy

OpenStudy (anonymous):

its too hard for me..please give the exact solution...thanks

OpenStudy (anonymous):

actually at this stage rather little leg work. solve in three steps 1) subtract x^2 from both sides 2) subtract 100 from both sides 3) divide both sides by 20

OpenStudy (anonymous):

i mean in terms of you have to actually write it down and solve it...can't just look at it anymore

OpenStudy (anonymous):

@leninyr i am sure it is not too hard for you now. follow the 3 steps and you will get it

OpenStudy (anonymous):

leinyr...do your own homework....he's helped you a ton

OpenStudy (anonymous):

oh yes of course! at some point you actually have to compute something. no getting around that.